The student government is selling raffle tickets to raise money for a school activity. A raffle ticket costs $1. There is 1 winning ticket out of the 2000 tickets sold. The winner gets a prize worth $75. Round your answers to the nearest cent.
What is the probability you win if you purchase one raffle ticket?

What is the expected value (to you) of one raffle ticket?

Answer by Boreal(15235)

1/2000.
It is $75*1/2000-$1*1999/2000=0.0375-0.9995=-$0.96.